A student in an organic chemistry lab is told to prepare C2H5Br by the reaction of:...

  1. A student in an organic chemistry lab is told to prepare C2H5Br by the reaction of:

C2H5OH(l) + PBr3(aq) → C2H5Br(l) + H3PO3(aq)

  1. What is the limiting reagent?
  2. What is the theoretical yield of C2H5Br?
  3. What is the mass of excess reagent remaining?
  4. IF the student actually obtains 36.0g of C2H5Br, what is the percent yield?
Answer #1

balanced reaction is

PBr3 + 3C2H5OH → 3C2H5Br + H3PO3

a. since 1 mol of PBr3 react with 3 mole of C2H5OH hence hence PBr3 is the limiting reagent

b. one mol of PBr3 give 3 mole of C2H5Br

mass = moles * Molar mass = 3 * 108.97 = 326.91 g

c. moles of excess reagent 3mol -1 mol = 2 mol

mass of excess reagent = 2 * 46 = 92 g

d. Perccent yield = (observed mass / theoretical mass ) * 100

= ( 36g / 326.91 ) *100 = 11%

PLEASE LEAVE THE POSITIVE FEEDBACK IF YOU LIKE THE SOLUTION